### v1

# Friedrich Pillichshammer, Kosuke Suzuki

Abstract
The discrepancy of a point set quantifies how well the points are distributed, with low-discrepancy point sets demonstrating exceptional uniform distribution properties. Such sets are integral to quasi-Monte Carlo methods, which approximate integrals over the unit cube for integrands of bounded variation. In contrast, quasi-uniform point sets are characterized by optimal separation and covering radii, making them well-suited for applications such as radial basis function approximation. This paper explores the quasi-uniformity properties of quasi-Monte Carlo point sets constructed from lattices. Specifically, we analyze rank-1 lattice point sets, Fibonacci lattice point sets, Frolov point sets, and $(n\boldsymbol{\alpha})$-sequences, providing insights into their potential for use in applications that require both low-discrepancy and quasi-uniform distribution. As an example, we show that the $(n\boldsymbol{\alpha})$-sequence with $\alpha_{j}=2^{j/(d+1)}$ for $j\in\{1,2,\ldots,d\}$ is quasi-uniform and has low-discrepancy.
Keywords: quasi-uniformity, low-discrepancy, quasi-Monte Carlo, lattice point sets, $(n\boldsymbol{\alpha})$-sequences
AMS subject classifications: 11K36, 11K38, 11J71
1 Introduction
In this paper we study two commonly required properties of point sets and sequences. On the one hand this is the property of quasi-uniformity, on the other hand the property of uniform distribution modulo 1 and low-discrepancy. Although the name would suggest that these properties may be related in some way, these two properties are fundamentally different as we will explain in Section 1.3. Both properties are well motivated and justified by important applications in numerical analysis, some of which will be indicated in Section 1.4. First we give the precise definitions of the two concepts under consideration.
1.1 Quasi-Uniformity
Let $\Omega\subseteq\mathbb{R}^{d}$ be a compact subset with $\mathrm{vol}(\Omega)>0$. Let $P$ be a point set in $\Omega$. For $p\in[1,\infty]$ let $\|\cdot\|_{p}$ denote the $\ell_{p}$ norm, i.e., $\|\boldsymbol{x}\|_{p}=(|x_{1}|^{p}+\cdots+|x_{d}|^{p})^{1/p}$ for $\boldsymbol{x}=(x_{1},\ldots,x_{d})\in\mathbb{R}^{d}$, with the obvious modifications for $p=\infty$.
Define the covering radius (in $\ell_{p}$ norm) by
$h_{p}(P;\Omega):=\sup_{\boldsymbol{x}\in\Omega}\min_{\boldsymbol{y}\in P}\|\boldsymbol{x}-\boldsymbol{y}\|_{p}.$
The covering radius can be described in the following way: If one places closed $\ell_{p}$ balls of radius $r$ around each point of $P$, then the smallest value of $r$ such that the union of the balls cover $\Omega$ is called the covering radius. In the literature, the covering radius is also known as dispersion (see, e.g., [28, Definition 6.2]), as fill distance (see, e.g., [20]), as mesh norm, or as minimax-distance criterion (see [34]).
The separation radius (in $\ell_{p}$ norm) is given by
$q_{p}(P;\Omega):=\min_{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in P\\ \boldsymbol{x}\neq\boldsymbol{y}\end{subarray}}\frac{\|\boldsymbol{x}-\boldsymbol{y}\|_{p}}{2}.$
The separation radius can be described in the following way: If one places open $\ell_{p}$ balls of radius $r$ around each point of $P$, then the separation radius is the largest value of $r$ such that none of these balls intersect in $\Omega$. In the literature, the separation radius is also known as packing radius or maximin-distance criterion; see [34].
The mesh ratio (in $\ell_{p}$ norm) is given by
$\rho_{p}(P;\Omega)=\frac{h_{p}(P;\Omega)}{q_{p}(P;\Omega)}.$
Since we are mainly dealing with the domain $[0,1]^{d}$ we write $h_{p}(P),q_{p}(P),\rho_{p}(P)$ instead of $h_{p}(P;[0,1]^{d}),q_{p}(P;[0,1]^{d}),\rho_{p}(P;[0,1]^{d})$ for brevity.
It is obvious that $\rho_{p}(P;\Omega)\geq 1$, at least if $\Omega$ is convex. Furthermore, it is known that the optimal order of the separation radius and the covering radius is $N^{-1/d}$. In more detail, we have
$h_{p}(P,\Omega)\geq\frac{c_{p,d}}{N^{1/d}}\quad\mbox{ and }\quad q_{p}(P,\Omega)\leq\frac{c_{p,d}^{\prime}}{N^{1/d}},$
for quantities $c_{p,d},c_{p,d}^{\prime}>0$, where the lower bound on the separation radius is taken from [34, Lemma 2.1] and the upper bound on the covering radius follows by comparing the volume of the balls around each point in $P$ to the volume of $\Omega$. Hence the mesh ratio of a sequence of point sets with an increasing number of points is bounded by a constant independent of the number of points if and only if both the covering radii and the separation radii are of the optimal order. In fact, in [34] it is shown that for any $N\in\mathbb{N}$ with $N>1$ one can always find point sets with $N$ points with mesh ratio at most $2$.
Definition 1.1.
Let $(P_{i})_{i\geq 1}$ be a sequence of point sets with increasing sizes, i.e., $|P_{1}|<|P_{2}|<\dots.$ Then we call $(P_{i})_{i\geq 1}$ a weakly quasi-uniform sequence of point sets if there exists a constant $C_{p}>0$ such that the mesh ratio is bounded, i.e.
$\rho_{p}(P_{i})\leq C_{p}\quad\forall i\in\mathbb{N}.$
Moreover, for a weakly quasi-uniform sequence $(P_{i})_{i\geq 1}$, if there exists a constant $c>1$ such that $|P_{i}|\leq c|P_{i+1}|$ for all $i\in\mathbb{N}$, we call $(P_{i})_{i\geq 1}$ a quasi-uniform sequence of point sets.
In this definition, the point sets do not need to be extensible. As mentioned above, the covering radius has its lower bound of order $N^{-1/d}$, while the separation radius has its upper bound of order $N^{-1/d}$. For $(P_{i})_{i\geq 1}$ to be quasi-uniform, it is required that the point sets $P_{i}$ attain the optimal order of both the covering and separation radii simultaneously.
For an infinite sequence of points, (weak) quasi-uniformity is similarly defined as follows:
Definition 1.2.
Let $\mathcal{S}=(\boldsymbol{x}_{n})_{n\geq 0}$ be an infinite sequence of points. We denote the set of the first $i$ points of $\mathcal{S}$ by $P_{i}=\{\boldsymbol{x}_{0},\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{i-1}\}$. Then, we call $\mathcal{S}$ a quasi-uniform sequence if there is a constant $C^{\prime}_{p}>0$ such that
$\rho_{p}(P_{i})\leq C^{\prime}_{p},\quad\forall i\in\mathbb{N}.$
Assume that there exists an increasing sequence $1\leq i_{1}<i_{2}<i_{3}<\cdots$ such that there is a constant $C^{\prime\prime}_{p}>0$ such that
$\rho_{p}(P_{i_{k}})\leq C^{\prime\prime}_{p},\quad\forall k\in\mathbb{N}.$
Then, we call $\mathcal{S}$ a weakly quasi-uniform sequence.
[…]
The following result is from [7].
Lemma 1.3.
Let $\mathcal{S}=(\boldsymbol{x}_{n})_{n\geq 0}$ be a sequence in $[0,1]^{d}$. Let $P_{i}=\{\boldsymbol{x}_{0},\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{i-1}\}$ denote the set of the first $i$ points of $\mathcal{S}$. Let $1\leq i_{1}<i_{2}<i_{3}<\dots$ be an increasing sequence.
  1. 1.
    Assume that $\mathcal{S}$ is a weakly quasi-uniform sequence such that the mesh ratio $\rho_{p}(P_{i_{k}})\leq C^{\prime\prime}_{p}$ for all $k\in\mathbb{N}$. Assume that there is a constant $c>1$ such that $i_{k+1}\leq ci_{k}$ for all $k\in\mathbb{N}$. Then $\mathcal{S}$ is a quasi-uniform sequence.
  2. 2.
    Assume that the covering radius $h_{p}(P_{i_{k}})\leq C_{p}i_{k}^{-1/d}$ for all $k\in\mathbb{N}$ for some constant $C_{p}>0$. Assume that there is a constant $c>1$ such that $i_{k+1}\leq ci_{k}$ for all $k\in\mathbb{N}$. Then there is a constant $D_{p}$ such that $h_{p}(P_{i})\leq D_{p}i^{-1/d}$ for all $i\in\mathbb{N}$.
  3. 3.
    Assume that the separation radius $q_{p}(P_{i_{k}})\geq C^{\prime}_{p}i_{k}^{-1/d}$ for all $k\in\mathbb{N}$ for some constant $C^{\prime}_{p}>0$. Assume that there is a constant $c>1$ such that $i_{k+1}\leq ci_{k}$ for all $k\in\mathbb{N}$. Then there is a constant $D^{\prime}_{p}$ such that $q_{p}(P_{i})\geq D^{\prime}_{p}i^{-1/d}$ for all $i\in\mathbb{N}$.
Brief summary of known results.
Quasi-uniform point sets play an important role for instance in the areas of design of computer experiments [13, 33, 36] and radial basis function approximation [37, 52].
The existence of quasi-uniform sequences with mesh ratio bounded by $2$ follows from [34] using a greedy algorithm which adds points one by one. Concrete explicit examples in dimension 1 are the famous van der Corput sequence in base $b$, which is quasi-uniform, or the Kronecker sequence ($(\{n\alpha\})_{n\geq 0}$, which is quasi-uniform if and only if $\alpha$ is a badly approximable number; see [16] and Section 3.3 of the present paper.
On the other hand, important sequences that seem at first sight to be candidates for quasi-uniform sequences do actually not possess this property. An example for such an instance is the Sobol’ sequence in dimension 2, which is not quasi-uniform, as shown in [17]. More examples can be found in [7], including Hammersley point sets, Faure sequence, and a class of Fibonacci polynomial lattices.
1.2 Uniform distribution modulo 1 and discrepancy
In 1916 Weyl [53] introduced the concept of uniform distribution modulo 1 of sequences. Accordingly, an infinite sequence $(\boldsymbol{x}_{n})_{n\geq 0}$ in $\mathbb{R}^{d}$ is called uniformly distributed modulo 1, if for every interval $[\boldsymbol{a},\boldsymbol{b})\subseteq[0,1)^{d}$ we have
$\lim_{N\rightarrow\infty}\frac{|\{n<N\ :\ \{\boldsymbol{x}_{n}\}\in[\boldsymbol{a},\boldsymbol{b})\}|}{N}={\rm Volume}([\boldsymbol{a},\boldsymbol{b})),$
[…]
A quantitative measure for how close the empirical distribution of a given point set in $[0,1)^{d}$ is to the uniform distribution is the notion of discrepancy. For a class $\mathcal{C}$ of measurable test sets in $[0,1)^{d}$ the discrepancy (with respect to $\mathcal{C}$) of an $N$-point set $P_{N}=\{\boldsymbol{x}_{0},\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N-1}\}$ in $[0,1)^{d}$ is defined as
$D^{\mathcal{C}}_{N}(P_{N})=\sup_{C\in\mathcal{C}}\left|\frac{|\{0\leq n<N\ :\ \boldsymbol{x}_{n}\in C\}|}{N}-{\rm Volume}(C)\right|.$
For an infinite sequence $\mathcal{S}=(\boldsymbol{x}_{n})_{n\geq 1}$ the discrepancy $D_{N}^{\mathcal{C}}(\mathcal{S})$ is defined as the discrepancy $D_{N}^{\mathcal{C}}(P_{N})$ of the initial segment $P_{N}=\{\boldsymbol{x}_{0},\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N-1}\}$ of $\mathcal{S}$.
Popular notions of discrepancy are the so-called star-discrepancy $D_{N}^{*}$, which refers to the class of all intervals of the form $[\boldsymbol{0},\boldsymbol{a})$ with $\boldsymbol{a}\in[0,1]^{d}$ as test sets, or the isotropic discrepancy $J_{N}$, where the class of test sets is given by all convex subsets $C$ of $[0,1)^{d}$. Those two notions are related in the following way: For every dimension $d\in\mathbb{N}$ there exists a quantity $C_{d}>0$ such that for every $N$-point set $P_{N}$ in $[0,1)^{d}$ we have
$D_{N}^{*}(P_{N})\leq J_{N}(P_{N})\leq C_{d}(D_{N}^{\ast}(P_{N}))^{1/d},$
see, e.g., [23, Chapter 2, Section 1, Theorem 1.6]. Furthermore, it is well-known that a sequence $\mathcal{S}$ in $[0,1)^{d}$ is uniformly distributed if and only if its star-discrepancy $D_{N}^{\ast}(\mathcal{S})$ tends to 0 for $N$ tending to infinity; see, e.g., [26, Theorem 2.15]. The same holds true with the star-discrepancy replaced by isotropic discrepancy.
With some abuse of nomenclature, we call a sequence $(P_{N_{i}})_{i\geq 1}$ of $N_{i}$-element point sets $P_{N_{i}}$ in $[0,1)^{d}$ with $N_{1}<N_{2}<N_{3}<\dots$ uniformly distributed modulo 1, if
$\lim_{i\rightarrow\infty}D_{N_{i}}^{*}(P_{N_{i}})=0.$
Brief summary of known results.
The following asymptotic results about discrepancies are known:
  • Star-discrepancy: For every $N$-point set $P_{N}$ in $[0,1)$ we have $D_{N}^{\ast}(P_{N})\geq 1/(2N)$. For every $d\in\mathbb{N}\setminus\{1\}$ there exist quantities $c_{d}>0$ and $\delta_{d}>0$ such that for every $N$-point set $P_{N}$ in $[0,1)^{d}$ we have
    $D_{N}^{\ast}(P_{N})\geq c_{d}\frac{(\log N)^{\frac{d-1}{2}+\delta_{d}}}{N},$
    where $\delta_{2}=1/2$ and $\lim_{d\rightarrow\infty}\delta_{d}=0$. This result, which is an important improvement of Roth’s lower bound [35], was shown in [4]. The result for $d=2$ is according to [40]. On the other hand, for every $N\in\mathbb{N}\setminus\{1\}$ there exists an $N$-point set $P_{N}$ in $[0,1)^{d}$, such that
    $D_{N}^{\ast}(P_{N})\ll_{d}\frac{(\log N)^{d-1}}{N}.$
    Examples are constructions according to Hammersley, Faure, Niederreiter, and others. See, e.g., [10] for an overview. It is still an open question whether the order of magnitude $(\log N)^{d-1}/N$ is best possible for the star-discrepancy of $N$-element point sets in $[0,1)^{d}$.
    It has become common to speak about “low-discrepancy” sequences or point sets if the corresponding star-discrepancy is of order of magnitude $(\log N)^{\kappa(d)}/N$ for a “small” (linearly in $d$) exponent $\kappa(d)$ or, slightly weaker, of order of magnitude $N^{-1+\delta}$ for every $\delta>0$. Low-discrepancy sequences and point sets are of utmost importance as underlying nodes of quasi-Monte Carlo integration rules; see, e.g., [10, 11, 23, 26, 28].
  • Isotropic discrepancy: In [41, Theorem 1] (see also [42, Theorem 13A]) Schmidt proved the following general lower bound for arbitrary point sets: For every dimension $d$ there exists a quantity $c_{d}>0$ such that for all $N$-element point sets $P_{N}$ in dimension $d$ we have
    $J_{N}(P_{N})\geq\frac{c_{d}}{N^{2/(d+1)}}.$
    This result is essentially (up to $\log$-factors) best possible as shown in [3] for $d=2$ and in [48] for $d\geq 3$ using probabilistic methods.
    For $N$-element lattice point sets $P_{N}$ in $[0,1)^{d}$ (which are the main focus of this work – for a formal definition see Section 2) we always have
    $J_{N}(P_{N})\geq\frac{c_{d}}{N^{1/d}},$
    for some $c_{d}>0$ depending only on $d$, as shown in [32].
    Upper bounds on the isotropic discrepancy of special point sets are given in [24, 25]. For example, $J_{N}(P_{N})\ll_{d}N^{-1/d}$ if $P_{N}$ is the $N$-element Hammersley net or the $N$-th initial segment of the Halton sequence in dimension $d$.
1.3 Quasi-uniformity versus uniform distribution
We now present examples that demonstrate the variety of scenarios that can arise concerning the quasi-uniformity and discrepancy of a point set. Here and throughout the paper we use the following notation: for functions $f,g:\mathbb{N}\rightarrow\mathbb{R}$ we write $f(N)\ll g(N)$ if there exists a quantity $C>0$ independent of $N$ such that $f(N)\leq C\ g(N)$ for all $N\in\mathbb{N}$. If we want to stress that $C$ depends on some parameters, say $d,p$, then this is indicated by writing $f(N)\ll_{d,p}g(N)$. If we have simultaneously $f(N)\ll g(N)$ and $g(N)\ll f(N)$ (of course with different involved quantities $C$), then we write $f(N)\asymp g(N)$ or, if a dependence on other parameters, say $d,p$, should be stressed, $f(N)\asymp_{d,p}g(N)$.
Quasi-uniformity does not imply uniform distribution.
Although the name quasi-uniform seems to suggest that quasi-uniformity is related to uniform distribution, it is well understood that this is actually not the case in general. For example, already in dimension $d=1$ the point sets
$\left\{0,\frac{1}{2N},\ldots,\frac{N-1}{2N}\right\}\cup\left\{\frac{1}{2},\frac{1}{2}+\frac{1}{4N},\ldots,\frac{1}{2}+\frac{2N}{4N}=1\right\}$
for $N\in\mathbb{N}$, are quasi-uniform, but are not uniformly distributed, since the star-discrepancy of this point set does not converge to 00 as $N\to\infty$.
A sequence of uniformly distributed lattice point sets which are not quasi-uniform.
Let $d,m\in\mathbb{N}$, $d>1$ and let $n_{1}=\cdots=n_{d-1}=m$ and $n_{d}=m^{2}$. Let $\boldsymbol{n}=\boldsymbol{n}(m)=(n_{1},n_{2},\ldots,n_{d})$ and
$\Gamma_{\boldsymbol{n}}=\left\{\left(\frac{k_{1}}{n_{1}},\ldots,\frac{k_{d}}{n_{d}}\right):0\leq k_{j}<n_{j},\ j=1,2,\ldots,d\right\},$
be the uniform grid with mesh sizes $n_{1},\ldots,n_{d}$. The total number of points $N=|\Gamma_{\boldsymbol{n}}|=\prod_{j=1}^{d}n_{j}=m^{d+1}$. Consider a box $B_{\boldsymbol{t}}=\prod_{j=1}^{d}[0,t_{j})$ for some $\boldsymbol{t}=(t_{1},\ldots,t_{d})\in(0,1]^{d}$. Let $k_{j}\in\mathbb{N}$ be the unique integer such that $k_{j}/n_{j}<t_{j}\leq(k_{j}+1)/n_{j}$, i.e., $k_{j}=\lceil t_{j}n_{j}\rceil-1$. Then the number of points from $\Gamma_{\boldsymbol{n}}$ that belong to $B_{\boldsymbol{t}}$ is $k_{1}k_{2}\cdots k_{d}$ and
$\lim_{m\to\infty}\frac{|B_{\boldsymbol{t}}\cap\Gamma_{\boldsymbol{n}}|}{N}=\lim_{m\to\infty}\prod_{j=1}^{d}\frac{\lceil t_{j}n_{j}\rceil-1}{n_{j}}=\prod_{j=1}^{d}t_{j}=\mathrm{Volume}(B_{\boldsymbol{t}}).$
Hence, the sequence of point sets $(\Gamma_{\boldsymbol{n}(m)})_{m\geq 1}$ is uniformly distributed.
On the other hand, the separation radius is
$q_{p}(\Gamma_{\boldsymbol{n}(m)})=\frac{1}{2n_{d}}=\frac{1}{2m^{2}}=\frac{1}{2N^{\frac{2}{d+1}}}=\frac{1}{2N^{\frac{1}{d}+\frac{d-1}{d(d+1)}}}.$
Since the separation is not of the optimal order (which would be $N^{-1/d}$), it follows that $(\Gamma_{\boldsymbol{n}(m)})_{m\geq 1}$ is not quasi-uniform.
A quasi-uniform sequence of lattice point sets is always uniformly distributed.
Lattice point sets are the major object of interest in this paper. The formal definition will be presented in Section 2. For the time being the following suffices: For $i\in\mathbb{N}$ let $T_{i}\in\mathbb{R}^{d\times d}$ and
$\mathbb{X}_{i}=T_{i}(\mathbb{Z}^{d})=\{T_{i}\boldsymbol{k}\ :\ \boldsymbol{k}\in\mathbb{Z}^{d}\}$
and $P_{N_{i}}=\mathbb{X}_{i}\cap[0,1)^{d}$ with $N_{i}=|P_{i}|$ and $N_{1}<N_{2}<\dots$. Assume that the mesh ratio $\rho_{p}(P_{N_{i}})\leq C<\infty$ for some constant $C$ independent of $N_{i}$. Then, necessarily, $q_{p}(P_{N_{i}})\ll N_{i}^{-1/d}$. From Lemma 2.9 (see Section 2 below) we know that for lattice point sets in $[0,1)^{d}$ the separation radius is of the same order of magnitude as the isotropic discrepancy. Hence we have
$D_{N_{i}}^{*}(P_{N_{i}})\leq J_{N_{i}}(P_{N_{i}})\asymp h_{p}(P_{N_{i}})\ll q_{p}(P_{N_{i}})\ll\frac{1}{N_{i}^{1/d}}.$ (1)
This implies that $(P_{N_{i}})_{i\geq 1}$ is uniformly distributed modulo 1.
A quasi-uniform sequence of lattice point sets with slowly converging discrepancy.
[…]
$D_{N}^{\ast}(\Lambda_{m})\asymp\frac{1}{m}=\frac{1}{N^{1/d}},$
where the number of points $|\Lambda_{m}|=N=m^{d}$; see, for instance, [10, Remark 3.33].
1.4 Motivation and Applications
The mesh-ratio and the related quantities covering radius and separation radius have important practical applications in a number of areas. One such area is experimental designs, including those for computer experiments. Among the many criteria used to evaluate the quality of experimental designs, one crucial criterion is the space-filling property; see, e.g., [13, 33, 36], which means one wants to have point sets with small covering radius. Recent theoretical advancements in scattered data approximation—particularly in Gaussian process regression—have shown that quasi-uniform point sets and sequences (i.e. bounded mesh ratio) are effective as experimental nodes, as they help reduce the predicted mean error, see [38, 50, 52, 54] among others. Quasi-uniform designs play also a role in statistical applications of experimental designs [20]. Another area is radial basis function approximation where the points are used as centers of the radial basis functions. In order to cover the entire space one wants the points to have a small covering radius, and in order to avoid numerical problems, one wants the points also to be well separated, see [44, 38, 52].
In the context of quasi-Monte Carlo methods, Sobol’ and Shukhman [46] studied the minimum distance (separation radius) for the 2-dimensional Sobol’ sequence. The first papers on maximized minimum distance in the context of quasi-Monte Carlo rules are [18] and [19], who study $2$ dimensional digital nets with maximized minimum distance for computer graphics applications. The dispersion (covering radius) of quasi-Monte Carlo point sets and applications thereof (quasi-Monte Carlo methods for optimization) is discussed in [28, Chapter 6].
On the other hand, the need of uniformly distributed sequences and point sets and sequences with low discrepancy is well enough known and has its legitimation in the theory of quasi-Monte Carlo methods; see [9, 10, 26, 28].
In Section 1.3 we discussed that the two properties, quasi uniformity and low discrepancy, are not necessarily related. Nevertheless, there may be situations in applications where the presence of both properties is advantageous. We describe one such situation in the following.
Let $K:[0,1]^{d}\times[0,1]^{d}\to\mathbb{R}$ be a continuous reproducing kernel and let $\mathcal{H}_{K}$ be the corresponding reproducing kernel Hilbert space (see [2] for information about the theory of reproducing kernel Hilbert spaces). Let $f\in\mathcal{H}_{K}$ and consider the problem of approximating the function $f$ given only function values. One method of approximating $f$ is to use kernel interpolation, that is, for a given point set $P_{N}=\{\boldsymbol{x}_{0},\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N-1}\}\subseteq[0,1]^{d}$, we use
$f(\boldsymbol{x})\approx\mathcal{A}(f)(\boldsymbol{x}):=\sum_{n=0}^{N-1}c_{n}K(\boldsymbol{x},\boldsymbol{x}_{n})$
with coefficients $c_{0},c_{1},\ldots,c_{N-1}$ that are determined in the following way: By demanding that $f(\boldsymbol{x}_{m})=\mathcal{A}(f)(\boldsymbol{x}_{m})$ for all $m\in\{0,1,\ldots,N-1\}$, we obtain the linear system
$f(\boldsymbol{x}_{m})=\sum_{n=0}^{N-1}c_{n}K(\boldsymbol{x}_{m},\boldsymbol{x}_{n})$
in the unknowns $c_{0},c_{1},\ldots,c_{N-1}$. We can write this in matrix form as
$\boldsymbol{f}=\mathcal{K}\boldsymbol{c},$ (2)
where $\boldsymbol{f}=(f(\boldsymbol{x}_{0}),f(\boldsymbol{x}_{1}),\ldots,f(\boldsymbol{x}_{N-1}))^{\top}$, $\mathcal{K}=[K(\boldsymbol{x}_{n},\boldsymbol{x}_{m})]_{n,m=0}^{N-1}$, and $\boldsymbol{c}=(c_{0},c_{1},\ldots,c_{N-1})^{\top}$.
In order to obtain a small approximation error, one approach is to use a point set $P_{N}$ as nodes of the involved integration rule for which the numerical integration error in some function space of is small (see [55, Section 3.2.1]). Depending on the function space, the integration error can often be related to the discrepancy of the underlying point set. Further, it is known from [28, Chapter 6] that many quasi-Monte Carlo point sets with small discrepancy have a covering radius which is asymptotically optimal.
Now consider the case where the point set $P_{N}$ has two points $\boldsymbol{x}_{i},\boldsymbol{x}_{j}$, which are very close to each other. Then the corresponding rows of the matrix $\mathcal{K}$ are almost the same and the linear system (2) is ill-conditioned. Hence in this situation it is advantageous to use low discrepancy point sets or point sets with small worst-case integration error in a suitable function space, which are also quasi-uniform.
1.5 Organization of the paper
The primal object of interest in this work are lattice point sets and, closely related, $(n\boldsymbol{\alpha})$-sequences. In Section 2 we recall the basic definition of lattices and lattice point sets and provide some tools and lemmas regarding lattices and mesh ratio.
The main results of this work will be presented in Section 3. In Section 3.1, Theorem 3.1, we present a construction of quasi-uniform nested shrinked and shifted lattices of low star-discrepancy by means of admissible lattices. Section 3.2 deals with rank-1 lattice point sets. In Corollary 3.3 we present an explicit construction of a rank-1 lattice point set in dimension two with bounded mesh ratio and low star-discrepancy. In Theorem 3.7 we consider the $d$-dimensional case and provide an existence result for rank-1 lattice point sets with bounded mesh ratio and low star-discrepancy Finally, in Section 3.3 we consider $(n\boldsymbol{\alpha})$-sequences and characterize exactly those $\boldsymbol{\alpha}\in\mathbb{R}^{d}$ for which the corresponding sequence is quasi-uniform; see Theorem 3.9. In a last step, we show that some of them even have very low star-discrepancy.
2 Background on lattices
In this work we put the main focus on lattice point sets and on $(n\boldsymbol{\alpha})$-sequences.
Definition 2.1 (Lattices).
Let $T\in\mathbb{R}^{d\times d}$ be an invertible matrix. This matrix generates a lattice in $\mathbb{R}^{d}$ given by
$\mathbb{X}=T(\mathbb{Z}^{d})=\{T\boldsymbol{k}\ :\ \boldsymbol{k}\in\mathbb{Z}^{d}\}.$
For later use we introduce also the dual lattice of $\mathbb{X}$ which is
$\mathbb{X}^{\bot}=\{\boldsymbol{h}\in\mathbb{Z}^{d}\ :\ \boldsymbol{h}\cdot\boldsymbol{y}\in\mathbb{Z}\mbox{ for all }\boldsymbol{y}\in\mathbb{X}\},$
where $\cdot$ denotes the standard inner product on the $\mathbb{R}^{d}$.
The separation radius of a lattice $\mathbb{X}$ over $\mathbb{R}^{d}$ satisfies
$q_{p}(\mathbb{X};\mathbb{R}^{d})=\min_{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in\mathbb{X}\\ \boldsymbol{x}\neq\boldsymbol{y}\end{subarray}}\dfrac{\|\boldsymbol{x}-\boldsymbol{y}\|_{p}}{2}=\min_{\boldsymbol{x}\in\mathbb{X}\setminus\{\boldsymbol{0}\}}\dfrac{\|\boldsymbol{x}\|_{p}}{2}$
and we define the covering radius by $h_{p}(\mathbb{X};\mathbb{R}^{d})=\sup_{\boldsymbol{x}\in\mathbb{R}^{d}}\min_{\boldsymbol{y}\in\mathbb{X}}\|\boldsymbol{x}-\boldsymbol{y}\|_{p}$.
Point sets with a lattice structure are often used as quadrature nodes in quasi-Monte Carlo algorithms; see, e.g., [9, 28, 45]. Due to the repeating structure, lattices are also the leading candidates for point sets with an optimal mesh ratio. Before studying the mesh ratio of some lattices, which are also of use in quasi-Monte Carlo algorithms, we first provide some relevant background material which will be used at later stages.
Successive minima.
The covering radius, separation radius and mesh ratio of a lattice are closely related to the notion of successive minima. For $i\in\{1,\ldots,d\}$, the $i$-th successive minima $\lambda_{i}^{(p)}$ (in $\ell_{p}$-norm) of a lattice $\mathbb{X}$ is defined as the minimum of all positive reals $\lambda$ such that $\mathbb{X}$ contains at least $i$ linearly independent vectors whose $\ell_{p}$-norms are at most $\lambda$. A vector which attains $\lambda_{1}^{(p)}$ is called a shortest vector (in $\ell_{p}$-norm). By definition, we have
$q_{p}(\mathbb{X};\mathbb{R}^{d})=\frac{\lambda_{1}^{(p)}}{2}.$ (3)
Furthermore, we have a bound on the covering radius as follows.
Lemma 2.2.
For a lattice $\mathbb{X}\subseteq\mathbb{R}^{d}$, we have
$h_{p}(\mathbb{X};\mathbb{R}^{d})\leq\dfrac{1}{2}\sum_{j=1}^{d}\lambda_{j}^{(p)}\leq\dfrac{d}{2}\lambda_{d}^{(p)}.$ (4)
Proof.
By the definition of the successive minima, we have linearly independent vectors $v_{1},\dots,v_{d}\in\mathbb{X}$ with $\|v_{j}\|_{p}\leq\lambda_{j}^{(p)}$ for $j\in\{1,2,\ldots,d\}$. Note that the lattice generated by these vectors is generally a sublattice of $\mathbb{X}$.
Let $\boldsymbol{x}\in\mathbb{R}^{d}$ and let $c_{1},\dots,c_{d}\in\mathbb{R}$ with $\boldsymbol{x}=\sum_{j=1}^{d}c_{j}v_{j}$. Let $c^{\prime}_{j}$ be the nearest integer to $c_{j}$ and let $\boldsymbol{y}:=\sum_{j=1}^{d}c^{\prime}_{j}v_{j}\in\mathbb{X}$. Then we have
$\|\boldsymbol{x}-\boldsymbol{y}\|_{p}=\left\|\sum_{j=1}^{d}(c^{\prime}_{j}-c_{j})v_{j}\right\|_{p}\leq\sum_{j=1}^{d}|c^{\prime}_{j}-c_{j}|\|v_{j}\|_{p}\leq\frac{1}{2}\sum_{j=1}^{d}\lambda_{j}^{(p)}.$
Thus, the first inequality of this lemma follows. The second inequality is obvious since we have $\lambda_{1}^{(p)}\leq\dots\leq\lambda_{d}^{(p)}$. ∎
We now show that the mesh ratio of the lattices from a family of lattices with relatively large shortest vectors is bounded by a constant, concluding that such lattices are quasi-uniform.
Lemma 2.3.
Let $\mathbb{X}=T(\mathbb{Z}^{d})$ be a lattice such that $\lambda_{1}^{(\infty)}\geq C({\rm det}\,T)^{1/d}$ for some $C>0$. Then we have
$\lambda_{d}^{(\infty)}\leq\frac{({\rm det}\,T)^{1/d}}{C^{d-1}}.$
Moreover, it follows that
$\rho_{\infty}(\mathbb{X};\mathbb{R}^{d})\leq\dfrac{d}{C^{d}}.$
Proof.
Minkowski’s second theorem (see, for example, [43, Theorem 16]), states that
$\prod_{j=1}^{d}\lambda_{j}^{(\infty)}\leq{\rm det}\,T.$
Combining this with the inequality $C({\rm det}\,T)^{1/d}\leq\lambda_{1}^{(\infty)}\leq\lambda_{j}^{(\infty)}$ for $j\in\{2,\dots,d-1\}$, we obtain the first claim. The second claim follows from (3), (4) and the first claim as
$\rho_{\infty}(\mathbb{X};\mathbb{R}^{d})=\dfrac{h_{\infty}(\mathbb{X};\mathbb{R}^{d})}{q_{\infty}(\mathbb{X};\mathbb{R}^{d})}\leq\dfrac{d\;\lambda_{d}^{(\infty)}/2}{\lambda_{1}^{(\infty)}/2}\leq\dfrac{d}{C^{d}}.\qed$
Lattice points inside the hypercube.
Let us now turn to the quasi-uniformity of lattice points in the unit hypercube. Specifically, for a $d$-dimensional lattice $\mathbb{X}=T(\mathbb{Z}^{d})$, we consider the point set $P:=\mathbb{X}\cap[0,1)^{d}$ and call $P=P(\mathbb{X})$ a lattice point set (in $[0,1)^{d}$). In general, this restriction changes the covering radius and the separation radius, but the impact is relatively small as the following lemma shows.
Lemma 2.4.
Let $\mathbb{X}=T(\mathbb{Z}^{d})$ be a $d$-dimensional lattice and $P=\mathbb{X}\cap[0,1)^{d}$. Assume that $|P|\geq 2$. Then, for any $p\in[1,\infty]$, we have
$q_{p}(P)\geq q_{p}(\mathbb{X};\mathbb{R}^{d})\qquad\mbox{ and }\qquad h_{\infty}(P)\leq 2h_{\infty}(\mathbb{X};\mathbb{R}^{d}).$ (5)
In particular,
$\rho_{\infty}(P)=\dfrac{h_{\infty}(P)}{q_{\infty}(P)}\leq 2\dfrac{h_{\infty}(\mathbb{X};\mathbb{R}^{d})}{q_{\infty}(\mathbb{X};\mathbb{R}^{d})}=2\rho_{\infty}(\mathbb{X};\mathbb{R}^{d}).$ (6)
Proof.
The result for the separation radius follows from its definition, since
$q_{p}(P)=\min_{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in P\\ \boldsymbol{x}\neq\boldsymbol{y}\end{subarray}}\dfrac{\|\boldsymbol{x}-\boldsymbol{y}\|_{p}}{2}\geq\min_{\begin{subarray}{c}\boldsymbol{x},\boldsymbol{y}\in\mathbb{X}\\ \boldsymbol{x}\neq\boldsymbol{y}\end{subarray}}\dfrac{\|\boldsymbol{x}-\boldsymbol{y}\|_{p}}{2}=q_{p}(\mathbb{X};\mathbb{R}^{d}).$
Let us move on to the covering radius. If $2h_{\infty}(\mathbb{X};\mathbb{R}^{d})\geq 1$, then the result is clear, since $\boldsymbol{0}\in P$ and hence
$h_{\infty}(P)\leq\sup_{\boldsymbol{x}\in[0,1]^{d}}\|\boldsymbol{x}\|_{\infty}=1\leq 2h_{\infty}(\mathbb{X};\mathbb{R}^{d}).$
Next, consider the case $12h_{\infty}(\mathbb{X};\mathbb{R}^{d})<$ 1. For $\boldsymbol{x}=(x_{1},\dots,x_{d})\in[0,1)^{d}$ put $x_{j}^{*}:=\max(0,x_{j}-2h_{\infty}(\mathbb{X};\mathbb{R}^{d}))$ for $j\in\{1,\dots,d\}$ and define the hypercube $D$ in $[0,1)^{d}$ by
$D:=\prod_{j=1}^{d}[x_{j}^{*},x_{j}^{*}+2h_{\infty}(\mathbb{X};\mathbb{R}^{d})].$
Let $\boldsymbol{z}$ be the center of $D$. By the definition of the covering radius, there exists $\boldsymbol{y}^{*}\in\mathbb{X}$ such that
$\|\boldsymbol{y}^{*}-\boldsymbol{z}\|_{\infty}\leq h_{\infty}(\mathbb{X};\mathbb{R}^{d}).$
This implies $\boldsymbol{y}^{*}\in D$ and thus $\boldsymbol{y}^{*}\in P$. Hence we have
$\min_{\boldsymbol{y}\in P}\|\boldsymbol{x}-\boldsymbol{y}\|_{\infty}\leq\|\boldsymbol{x}-\boldsymbol{y}^{*}\|_{\infty}\leq 2h_{\infty}(\mathbb{X};\mathbb{R}^{d}).$
Since this bound is independent of the location of $\boldsymbol{x}\in[0,1)^{d}$, we obtain the desired result.
The result (6) is clear. ∎
Remark 2.5.
Although we omit the details, the result of the lemma holds more generally for any hypercube of the form $\Omega=[a_{1},b_{1})\times\cdots\times[a_{b},b_{d})$ with $a_{j}<b_{j}$ for all $j\in\{1,\ldots,d\}$. More precisely, with $P_{\Omega}=\mathbb{X}\cap\Omega$ such that $|P_{\Omega}|\geq 2$, for any $p\in[1,\infty]$ we have
$q_{p}(P_{\Omega};\Omega)\geq q_{p}(\mathbb{X};\mathbb{R}^{d})\qquad\text{ and }\qquad h_{\infty}(P_{\Omega};\Omega)\leq 2h_{\infty}(\mathbb{X};\mathbb{R}^{d}).$
This implies that any shifted lattice in the unit hypercube, defined by
$P=\left\{T\boldsymbol{k}-\boldsymbol{\delta}\ :\ \boldsymbol{k}\in\mathbb{Z}^{d}\right\}\cap[0,1)^{d}\quad\text{with $\boldsymbol{\delta}\in\mathbb{R}^{d}$,}$
also satisfies the same bound (6) on the mesh ratio.
With (6), we can show that the “shrinking” strategy works to construct quasi-uniform point sets and sequences.
Lemma 2.6.
Let $\mathbb{X}=T(\mathbb{Z}^{d})$ be a $d$-dimensional Euclidean lattice. For $a>0$, define
$P_{a}:=a^{-1}\mathbb{X}\cap[0,1)^{d},$
and assume that $|P_{a}|\geq 2$. Then, it holds that
$\rho_{\infty}(P_{a})\leq 2\rho_{\infty}(\mathbb{X};\mathbb{R}^{d})\quad\mbox{for every }a>0\mbox{ such that }|P_{a}|\geq 2.$
In particular, for any sequence $0<a_{1}<a_{2}<\dots$ with $a_{i}\to\infty$ as $i\to\infty$, the sequence of point sets $(P_{a_{i}})_{i\geq 1}$ is weakly quasi-uniform over $[0,1)^{d}$.
Moreover, if the sequence $0<a_{1}<a_{2}<\dots$ is such that there is a constant $c>1$ such that $N_{a_{i+1}}\leq cN_{a_{i}}$ for all $i$, then the sequence $(P_{a_{i}})_{i\geq 1}$ is quasi-uniform.
Proof.
It is evident that $a^{-1}\mathbb{X}$ is a $d$-dimensional lattice and satisfies
$h_{p}(a^{-1}\mathbb{X};\mathbb{R}^{d})=a^{-1}h_{p}(\mathbb{X};\mathbb{R}^{d})\quad\text{and}\quad q_{p}(a^{-1}\mathbb{X};\mathbb{R}^{d})=a^{-1}q_{p}(\mathbb{X};\mathbb{R}^{d}),$ (7)
for all $p\in[1,\infty]$. From (6), it follows that
$\rho_{\infty}(P_{a})\leq 2\frac{h_{\infty}(a^{-1}\mathbb{X};\mathbb{R}^{d})}{q_{\infty}(a^{-1}\mathbb{X};\mathbb{R}^{d})}=2\frac{a^{-1}h_{\infty}(\mathbb{X};\mathbb{R}^{d})}{a^{-1}q_{\infty}(\mathbb{X};\mathbb{R}^{d})}=2\rho_{\infty}(\mathbb{X};\mathbb{R}^{d}).$
Since this bound is independent of $a$, the claim follows. ∎
Remark 2.7.
Let $\mathbb{X}=T(\mathbb{Z}^{d})$ be a lattice with $\rho_{p}(\mathbb{X};\mathbb{R}^{d})<\infty$. Let $P_{a_{i}}=a_{i}^{-1}\mathbb{X}\cap[0,1)^{d}$ and $N_{a_{i}}=|P_{a_{i}}|$. The covering radius for any finite point set $P$ satisfies $h_{\infty}(P)\gg|P|^{-1/d}$ and the separation radius satisfies $p_{\infty}(P)\ll|P|^{-1/d}$. Using this together with (5) and (7) we obtain
$N_{a_{i}}^{-1/d}\ll h_{\infty}(P_{a_{i}})\leq 2a_{i}^{-1}h_{\infty}(\mathbb{X}).$
This implies that $N_{a_{i}}\gg a_{i}^{d}$. Using the separation radius we similarly obtain $N_{a_{i}}\ll a_{i}^{d}$.
Let $z>1$ be a real number and $a_{0}>0$. Let $a_{i}=a_{0}z^{i}$ for all $i\in\mathbb{N}$. Then $N_{a_{i+1}}\ll a_{i+1}^{d}=a_{0}^{d}z^{id}z^{d}\ll z^{d}N_{a_{i}}$ for all $i$. Hence for $a_{i}=a_{0}z^{i}$ there is a $c>1$ such that $N_{a_{i+1}}\leq cN_{a_{i}}$. Under the assumptions from Lemma 2.6, with this choice of $a_{i}$, we obtain a quasi-uniform sequence of point sets.
Remark 2.8.
Let $n>1$ be an integer, and consider the case where $a_{i}=a_{0}n^{i}$ for all $i\in\mathbb{N}$ with some $a_{0}>0$. We see that
$a_{i+1}^{-1}\mathbb{X}$ =$=\left\{a_{0}^{-1}n^{-i-1}T\boldsymbol{k}\ :\ \boldsymbol{k}\in\mathbb{Z}^{d}\right\}$
=$=\left\{a_{0}^{-1}n^{-i-1}T\boldsymbol{k}\ :\ \boldsymbol{k}\in(n\mathbb{Z})^{d}\right\}\cup\left\{a_{0}^{-1}n^{-i-1}T\boldsymbol{k}\ :\ \boldsymbol{k}\in\mathbb{Z}^{d}\setminus(n\mathbb{Z})^{d}\right\}$
=$=\left\{a_{0}^{-1}n^{-i}T\boldsymbol{k}\ :\ \boldsymbol{k}\in\mathbb{Z}^{d}\right\}\cup\left\{a_{0}^{-1}n^{-i-1}T\boldsymbol{k}\ :\ \boldsymbol{k}\in\mathbb{Z}^{d}\setminus(n\mathbb{Z})^{d}\right\}$
$=a_{i}^{-1}\mathbb{X}\cup\left\{a_{0}^{-1}n^{-i-1}T\boldsymbol{k}\ :\ \boldsymbol{k}\in\mathbb{Z}^{d}\setminus(n\mathbb{Z})^{d}\right\}\supseteq a_{i}^{-1}\mathbb{X}.$
Hence, the sequence $(a_{i}^{-1}\mathbb{X})_{i\geq 1}$ is a nested sequence (in quasi-Monte Carlo theory this property is referred to as “extensibility”) and therefore $(P_{a_{i}})_{i\geq 1}$, where $P_{a_{i}}=a_{i}^{-1}\mathbb{X}\cap[0,1)^{d}$, is also a nested sequence of quasi-uniform point sets.
The spectral test and enhanced trigonometric degree of lattices.
The following result can be easily brought together from well-known results in the literature. Let $\mathbb{X}=T(\mathbb{Z}^{d})$ be a $d$-dimensional lattice and $P=\mathbb{X}\cap[0,1)^{d}$ with $|P|=N$. The spectral-test of the lattice $\mathbb{X}$, is
$\sigma_{N}(P)=\frac{1}{\min_{\boldsymbol{h}\in\mathbb{X}^{\bot}\setminus\{\boldsymbol{0}\}}\|\boldsymbol{h}\|_{2}}$
and the enhanced trigonometric degree of the lattice $\mathbb{X}$ is
$\kappa_{N}(P)=\min_{\boldsymbol{h}\in\mathbb{X}^{\bot}\setminus\{\boldsymbol{0}\}}\|\boldsymbol{h}\|_{1}.$
Like discrepancy, the spectral-test and enhanced trigonometric degree are important figure-of-merits for lattices in the context of numerical integration. For general information we refer to, e.g., [9, Section 1.8].
Lemma 2.9.
Let $d\in\mathbb{N}$ and $\mathbb{X}=T(\mathbb{Z}^{d})$ be a $d$-dimensional lattice and $P=\mathbb{X}\cap[0,1)^{d}$ with $|P|=N$. Then for any $p\in[1,\infty]$ we have
$h_{p}(P)\asymp_{p,d}J_{N}(P)\asymp_{p,d}\sigma_{N}(P)\asymp_{p,d}\frac{1}{\kappa_{N}(P)},$
where $J_{N}(P)$ denotes the isotropic discrepancy of $P$, $\sigma_{N}(P)$ is the spectral-test of the lattice $\mathbb{X}$ and $\kappa_{N}(P)$ is the enhanced trigonometric degree of the lattice $\mathbb{X}$.
Proof.
The equivalence $h_{p}(P)\asymp_{p,d}J_{N}(P)$ is [47, Proposition 8] (with $\gamma=\infty$), the equivalence $J_{N}(P)\asymp_{p,d}\sigma_{N}(P)$ is [32, Theorem 2], and the equivalence $\sigma_{N}(P)\asymp_{p,d}1/\kappa_{N}(P)$ is clear from the definition and the equivalence of $\ell_{p}$-norms. ∎
In the next sections we provide examples of lattice constructions in the unit cube which have both, small discrepancy and bounded mesh ratio.
3 The main results
3.1 Admissible lattices - Frolov’s construction
Let $\mathbb{X}$ be an admissible lattice, that is, $\mathbb{X}$ satisfies
$\inf_{\boldsymbol{x}\in\mathbb{X}\setminus\{\boldsymbol{0}\}}\prod_{j=1}^{d}\left|x_{j}\right|>0.$
An example of such a lattice was proposed by Frolov [14] (see also [51]). The discrepancy of admissible lattices was studied in [15, 44].
Let $\mathbb{X}=T(\mathbb{Z}^{d})$ be an admissible lattice. From [44, Corollary 2.1, item 1] it is known that for any shift $\boldsymbol{\delta}\in\mathbb{R}^{d}$ and any $a>0$ the shrinked, shifted lattice $P$ inside the hypercube of the form
$P_{a,\boldsymbol{\delta}}=\left\{a^{-1}\left(T\boldsymbol{k}-\boldsymbol{\delta}\right)\ :\ \boldsymbol{k}\in\mathbb{Z}^{d}\right\}\cap[0,1)^{d}$ (8)
has star-discrepancy of order of magnitude $(\log N)^{d-1}/N$, where $N=|P_{a,\boldsymbol{\delta}}|$.
Using this result, Remark 2.7 and Remark 2.8, we can obtain a nested sequence of shifted lattice point sets with very low (possibly best possible) star-discrepancy, which is also quasi-uniform.
Theorem 3.1.
Let $\mathbb{X}=T(\mathbb{Z}^{d})$ be an admissible lattice. For $a>0$ and $\boldsymbol{\delta}\in\mathbb{R}^{d}$ let $P_{a,\boldsymbol{\delta}}$ be the point set from (8). For an integer $n\geq 2$, let $a_{i}=a_{0}n^{i}$ for all $i\in\mathbb{N}$ with some $a_{0}>0$ and let $\boldsymbol{\delta}_{i}\in\mathbb{R}^{d}$ for all $i\in\mathbb{N}$. Then the sequence $(P_{a_{i},\boldsymbol{\delta}_{i}})_{i\geq 1}$ is a quasi-uniform sequence of nested point sets whose star-discrepancy is of order of magnitude $(\log N_{i})^{d-1}/N_{i}$, where $N_{i}=|P_{a_{i},\boldsymbol{\delta}_{i}}|$ for all $i\in\mathbb{N}$.
For practical implementations of such lattices, see also [6, 21, 49].
3.2 Rank-1 lattice point sets
An important instance of lattice point sets are so-called rank-1 lattice point sets. Let $N\in\mathbb{N}$ and $\boldsymbol{g}=(g_{1},\ldots,g_{d})\in\{0,1,\ldots,N-1\}^{d}$ with $\gcd(g_{i},N)=1$ for all $i\in\{1,\ldots,d\}$. Let $\mathbb{X}(\boldsymbol{g},N)=\{T\boldsymbol{k}\ :\ \boldsymbol{k}\in\mathbb{Z}^{d}\}$, where, for $g_{1}\neq 0$,
$T=\begin{pmatrix}1/N&0&\ldots&\ldots&\ldots&0\\ g_{1}^{-1}g_{2}/N&1&0&\ldots&\ldots&0\\ \vdots&0&\ddots&\ddots&&\vdots\\ \vdots&\vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\vdots&&\ddots&\ddots&0\\ g_{1}^{-1}g_{d}/N&0&\ldots&\ldots&0&1\end{pmatrix},$
and for $g_{1}=0$,
$T=\begin{pmatrix}1&0&\ldots&\ldots&\ldots&0\\ g_{2}/N&1&0&\ldots&\ldots&0\\ \vdots&0&\ddots&\ddots&&\vdots\\ \vdots&\vdots&\ddots&\ddots&\ddots&\vdots\\ \vdots&\vdots&&\ddots&\ddots&0\\ g_{d}/N&0&\ldots&\ldots&0&1\end{pmatrix},$
be a so-called rank-1 lattice and $P(\boldsymbol{g},N)=\mathbb{X}(\boldsymbol{g},N)\cap[0,1)^{d}$, i.e.
$P(\boldsymbol{g},N)=\left\{\left\{\frac{k}{N}\boldsymbol{g}\right\}\ :\ k\in\{0,1,\ldots,N-1\}\right\}$
be the corresponding rank-1 lattice point set consisting of $N$ points in $[0,1)^{d}$. In the present case, the dual lattice is given by
$\mathbb{X}^{\bot}(\boldsymbol{g},N)=\{\boldsymbol{h}\in\mathbb{Z}^{d}\ :\ \boldsymbol{g}\cdot\boldsymbol{h}\equiv 0\pmod{N}\},$
see [9, Section 1.2]. Detailed expositions on rank-1 lattice point sets in the theory of quasi-Monte Carlo integration can be found in the books [9, 28, 45].
In dimension $2$ we can obtain explicit constructions of rank-1 lattice point sets with small discrepancy, which are quasi-uniform. However, in higher dimensions, we can only prove an existence result.
Rank-1 lattice point sets in dimension two.
We study the quasi-uniformity of rank-1 lattice point sets $P((1,g),N)\subseteq[0,1)^{2}$ for certain $g\in\{1,2,\ldots,N-1\}$. In order to study the covering and separation radii of such lattice point sets, we need to bound the successive minima; see Section 2. The candidates of a shortest vector of a planar lattice generated by an integer matrix are given by the continued fraction of $g/N$; see [12]. Considering all candidates, we can show a bound on the mesh ratio. For the sake of completeness, we provide a full proof. Before we do so, we provide some background on continued fraction expansions.
A continued fraction represents a positive real number $x$ in the form
$x=a_{0}+\cfrac{1}{a_{1}+\cfrac{1}{a_{2}+\cfrac{1}{a_{3}+\cfrac{1}{\ddots}}}}=:[a_{0};a_{1},a_{2},a_{3},\dots]$
where $a_{0},a_{1},a_{2},a_{3},\dots$ are positive integers. The $n$-th convergent of $x$ is given by $r_{n}=[a_{0};a_{1},\dots,a_{n}]=p_{n}/q_{n}$ with $\gcd(p_{n},q_{n})=1$, where $p_{n}$ and $q_{n}$ can be computed recursively
$p_{n}$ $=a_{n}p_{n-1}+p_{n-2},$
$q_{n}$ $=a_{n}q_{n-1}+q_{n-2},$
with initial conditions
$p_{-2}=0,\,p_{-1}=1,q_{-2}=1,\,q_{-1}=0.$
If $x$ is rational, the continued fraction expansion is finite and uniquely determined if we write it in the form
$x=[a_{0};a_{1},\dots,a_{\ell}]\qquad\text{with }a_{\ell}=1\,\text{ for }\ell\in\mathbb{N}.$ (9)
The convergents are good Diophantine approximations of $x$. That is, for any real $x$ the $n$-th convergent $r_{n}=p_{n}/q_{n}$ satisfies
$\frac{1}{q_{n}(q_{n}+q_{n+1})}\leq\left|x-\frac{p_{n}}{q_{n}}\right|\leq\frac{1}{q_{n}q_{n+1}}.$ (10)
Furthermore, the $n$-th convergent is a best rational approximation of $x$ in the sense that, for any integers $1\leq h<q_{n+1}$ and $b$ we have
$|hx-b|\geq|q_{n}x-p_{n}|.$ (11)
For more information on continued fraction expansions and their use in Diophantine approximation theory, we refer to [22].
Now we return to the mesh-ratio of rank-1 lattice point sets in dimension 2.
Theorem 3.2.
Let $g,N\in\mathbb{N}$ with $\gcd(g,N)=1$ and assume that $x=g/N$ has continued fraction expansion of the form (9). Let
$T=\begin{pmatrix}1/N&0\\ g/N&1\end{pmatrix}\qquad\mbox{ and }\qquad\mathbb{X}=\mathbb{X}((1,g),N)=T(\mathbb{Z}^{2}).$
Then, denoting $K=\max(a_{1},\dots,a_{\ell})$, we have
$\lambda_{1}^{(\infty)}\geq\dfrac{1}{\sqrt{(K+2)N}}.$
Furthermore, we have
$\rho_{\infty}(\mathbb{X};\mathbb{R}^{2})\leq 2(K+2).$
Proof.
To show the first claim, it suffices to show that $\|\boldsymbol{x}\|_{\infty}\geq 1/\sqrt{(K+2)N}$ holds true for any $\boldsymbol{x}\in\mathbb{X}\setminus\{\boldsymbol{0}\}$. Now every $\boldsymbol{x}\in\mathbb{X}\setminus\{\boldsymbol{0}\}$ can be written as $(t/N,tg/N-u)$ for some $t,u\in\mathbb{Z}$ with $(t,u)\neq(0,0)$. By symmetry, we only need to consider the case $t\geq 0$. We distinguish three cases:
  • If $t=0$, then $u\neq 0$ and thus we have $\|(t/N,tg/N-u)\|_{\infty}=|u|\geq 1\geq 1/\sqrt{(K+2)N}.$
  • If $t\geq N$, then we have $\|(t/N,tg/N-u)\|_{\infty}\geq t/N\geq 1\geq 1/\sqrt{(K+2)N}$.
  • Now let $0<t<N$. Let $p_{n}/q_{n}$ be the $n$-th convergent of $x=g/N$ for $n=0,\dots,\ell$. Take $j$ with $q_{j}\leq t<q_{j+1}$. Then it follows from (11) and (10) that
    $\left|\dfrac{tg}{N}-u\right|\geq\left|\dfrac{q_{j}g}{N}-p_{j}\right|=q_{j}\left|\dfrac{g}{N}-\dfrac{p_{j}}{q_{j}}\right|\geq\dfrac{1}{q_{j}+q_{j+1}}=\dfrac{1}{q_{j}+(a_{j+1}q_{j}+q_{j-1})}\geq\dfrac{1}{(K+2)q_{j}}.$
    Thus we have
    $\left\|\left(\dfrac{t}{N},\dfrac{tg}{N}-u\right)\right\|_{\infty}\geq\max\left(\dfrac{q_{j}}{N},\dfrac{1}{(K+2)q_{j}}\right)\geq\dfrac{1}{\sqrt{(K+2)N}},$
    where we use $\max(a,b)\geq\sqrt{ab}$ for $a,b>0$ in the last inequality.
These three cases imply the first claim. Noting that ${\rm det}\,T=1/N$, the second claim follows directly from the first one and Lemma 2.3. ∎
Combining Theorem 3.2 and (6), we obtain the following corollary:
Corollary 3.3.
Let $g,N\in\mathbb{N}$ with $\gcd(g,N)=1$ and $P=P((1,g),N)$ be the corresponding rank-1 lattice point set. Assume $x=g/N$ has continued fraction expansion of the form (9) and $K:=\max(a_{1},\dots,a_{\ell})$. Then we have
$\rho_{\infty}(P)\leq 4(K+2).$
Remark 3.4.
It is well-known that the discrepancy of $2$-dimensional rank-$1$ lattice point sets is also strongly related to the continued fraction expansion of $g/N$. Under the same assumption of Corollary 3.3, the star-discrepancy of $P((1,g),N)$ satisfies
$D_{N}^{*}(P((1,g),N))\ll K\frac{\log N}{N}$
with an absolute implied constant [28, (5.39) and the following discussion]. Thus such point sets are also low-discrepancy point sets.
Remark 3.5.
Let $F_{1}=F_{2}=1$ and $F_{k}=F_{k-1}+F_{k-2}$ for $k\geq 3$ be the Fibonacci numbers. Then the Fibonacci lattice point set is given by
$\mathcal{F}_{m}=\left\{\left(\frac{n}{F_{m}},\left\{\frac{nF_{m-1}}{F_{m}}\right\}\right)\ :\ n=0,1,2,\ldots,F_{m}-1\right\},$
i.e., in general notation, $\mathcal{F}_{m}=P((1,g),N)$ with $g=F_{m-1}$ and $N=F_{m}$.
Since it is well-known that $F_{m-1}/F_{m}=[0;1,1,\dots,1]$, it follows from Theorem 3.2 with $K=1$ and $F_{m+1}\leq 2F_{m}$ that the family of Fibonacci lattices is quasi-uniform and also of low-discrepancy. Since the explicit shortest basis of (the underlying planar lattice of) Fibonacci lattices is known (see [1, 29]), one can obtain the exact separation radius and tighter bounds on the covering radius and the mesh ratio of the Fibonacci lattices.
Remark 3.6.
The minimum distance of rank-1 lattice point sets has been studied in the context of the traveling salesman path (see [30, 31]). Consider the sequence of lattices $P_{k}=P((1,g_{k}),N_{k})$ where $g_{k}/N_{k}=[0;a_{1},\cdots,a_{2k+1}]$ with $a_{2i}=1$ and $a_{2i+1}=2$ holds. It is shown in [31] that the shape of the lattices $P_{k}$ converges to the hexagonal lattice and the minimum distance of $P_{k}$ converges to $\sqrt{2/(\sqrt{3}N)}$. This sequence is also of low-discrepancy and, according to Corollary 3.3, also quasi-uniform.
Rank-1 lattice point sets in arbitrary dimension $d$.
We now turn the case of general dimension $d\in\mathbb{N}$ and provide an existence result of rank-1 lattice point sets with bounded mesh ratio and low star-discrepancy.
According to Lemma 2.9, the covering radius of $P(\boldsymbol{g},N)$ satisfies
$h_{p}(P(\boldsymbol{g},N))\ll_{d,p}\sigma(P(\boldsymbol{g},N))=\frac{1}{\min_{\boldsymbol{h}\in\mathbb{X}^{\bot}(\boldsymbol{g},N)\setminus\{\boldsymbol{0}\}}\|\boldsymbol{h}\|_{2}},$
where $\sigma(P(\boldsymbol{g},N))$ is the spectral test of $\mathbb{X}(\boldsymbol{g},N)$. Furthermore, the separation distance is
$q_{p}(P(\boldsymbol{g},N))=\frac{1}{2}\min_{\boldsymbol{x}\in P(\boldsymbol{g},N)\setminus\{\boldsymbol{0}\}}\|\boldsymbol{x}\|_{p}\geq\frac{1}{2}\min_{\boldsymbol{x}\in\mathbb{X}(\boldsymbol{g},N)\setminus\{\boldsymbol{0}\}}\|\boldsymbol{x}\|_{p}.$
These considerations imply
$\rho_{p}(P(\boldsymbol{g},N))\ll_{d,p}\frac{1}{\min_{\boldsymbol{h}\in\mathbb{X}^{\bot}(\boldsymbol{g},N)\setminus\{\boldsymbol{0}\}}\|\boldsymbol{h}\|_{p}\min_{\boldsymbol{h}\in\mathbb{X}(\boldsymbol{g},N)\setminus\{\boldsymbol{0}\}}\|\boldsymbol{h}\|_{p}}.$
Write
$\kappa(\mathbb{X})=\min_{\boldsymbol{h}\in\mathbb{X}\setminus\{\boldsymbol{0}\}}\|\boldsymbol{h}\|_{1}.$
Then we have
$\rho_{p}(P(\boldsymbol{g},N))\ll_{d,p}\frac{1}{\kappa(\mathbb{X}(\boldsymbol{g},N))\kappa(\mathbb{X}^{\bot}(\boldsymbol{g},N))}.$ (12)
The following theorem shows the existence of rank-1 lattice point sets with bounded mesh ratio and low star-discrepancy.
Theorem 3.7.
For every dimension $d\in\mathbb{N}$, every $p\in[1,\infty]$ and every $\delta\in(0,1)$ there exist quantities $B_{d,p,\delta},C_{d,p,\delta}>0$, which only depend on $d$, $p$ and $\delta$, with the following property: for every prime number $N$ and every $\delta\in(0,1)$ there exist at least $\lceil\delta N^{d}\rceil$ vectors $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ such that the mesh ratio is bounded
$\rho_{p}(P(\boldsymbol{g},N))\leq B_{d,p,\delta},$
and the star-discrepancy satisfies
$D_{N}^{\ast}(P(\boldsymbol{g},N))\leq C_{d,p,\delta}\frac{(\log N)^{d}}{N}.$
Proof.
Let $\delta\in(0,1)$. It follows from a standard averaging argument, see, e.g., [9, Lemma 2.13 and Sec. 5.1], that there exists a $C^{\prime}_{d,\delta}>0$ with the property, that
$D_{N}^{\ast}(P(\boldsymbol{g},N))\leq C_{d,\delta}^{\prime}\frac{(\log N)^{d}}{N}$ (13)
for at least $((2+\delta)/3)N^{d}$ generating vectors $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$.
Now we consider the existence of rank-1 lattice point sets with bounded mesh ratio. From (12) it follows that we need prove bounds on $\kappa(\mathbb{X}(\boldsymbol{g},N))$ and $\kappa(\mathbb{X}^{\perp}(\boldsymbol{g},N))$.
First consider $\kappa(\mathbb{X}^{\bot}(\boldsymbol{g},N))$. The proof follows [9, Proof of Proposition 1.61], for completeness we include it here.
For a given $\boldsymbol{h}=(h_{1},\ldots,h_{d})\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}$ with $|h_{j}|<N$ for all $j\in\{1,\dots,d\}$, there are $N^{d-1}$ choices of $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ such that $\boldsymbol{h}\cdot\boldsymbol{g}\equiv 0\pmod{N}$. Furthermore
|$|\{\boldsymbol{h}\in\mathbb{Z}^{d}\ :\ \|\boldsymbol{h}\|_{1}=\ell\}|\leq 2^{d}\binom{\ell+d-1}{d-1}.$
Let $\kappa<N$ be a given positive integer (note that we always have $\kappa(\mathbb{X}^{\bot}(\boldsymbol{g},N))<N$). Put
$A_{\kappa}:=\{\boldsymbol{h}\in\mathbb{Z}^{d}\ :\ \|\boldsymbol{h}\|_{1}\leq\kappa\}.$
Then
$|A_{\kappa}|\leq 2^{d}\sum_{\ell=0}^{\kappa}\binom{\ell+d-1}{d-1}=2^{d}\binom{\kappa+d}{d}.$
For every $\boldsymbol{h}$ in the set $\{\boldsymbol{h}\in\mathbb{Z}^{d}\ :\ \|\boldsymbol{h}\|_{1}\leq\kappa\}$ there are $N^{d-1}$ choices $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ such that $\boldsymbol{h}\cdot\boldsymbol{g}\equiv 0\pmod{N}$. For every such $\boldsymbol{g}$ we have $\kappa(\mathbb{X}^{\bot}(\boldsymbol{g},N))\leq\kappa$. Therefore
|$|\{\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}\ :\ \kappa(\mathbb{X}^{\bot}(\boldsymbol{g},N))\leq\kappa\}|\leq N^{d-1}2^{d}\binom{\kappa+d}{d}.$
Note that the total number of possible generators $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ is $N^{d}$. Thus, if
$N^{d-1}2^{d}\binom{\kappa+d}{d}<\frac{1-\delta}{3}N^{d},$ (14)
then the number of $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ such that $\kappa(\mathbb{X}^{\bot}(\boldsymbol{g},N))>\kappa$ is at least $((2+\delta)/3)N^{d}$. We estimate
$2^{d}\binom{\kappa+d}{d}\leq\frac{2^{d}(\kappa+d)^{d}}{d!}.$ (15)
Thus (14) is satisfied if $32^{d}(\kappa+d)^{d}/d!<(1-\delta)N/$ 3, i.e., for $\kappa=\lceil 2^{-1}(d!\,(1-\delta)N/3)^{1/d}\rceil-d-1$. Hence, for at least $((2+\delta)/3)N^{d}$ many $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ we have
$\kappa(\mathbb{X}^{\bot}(\boldsymbol{g},N))\geq B_{d,\delta}^{\prime}N^{1/d}$ (16)
for some $B_{d,\delta}^{\prime}>0$, depending only on $d$ and $\delta$.
Now we consider $\kappa(\mathbb{X}(\boldsymbol{g},N))$. The elements in $\mathbb{X}(\boldsymbol{g},N)$ are of the form $(1/N)\boldsymbol{h}$ with $\boldsymbol{h}\in\mathbb{Z}^{d}$. Hence
$\kappa(\mathbb{X}(\boldsymbol{g},N))=\frac{1}{N}\min_{(1/N)\boldsymbol{h}\in\mathbb{X}(\boldsymbol{g},N)\setminus\{\boldsymbol{0}\}}\|\boldsymbol{h}\|_{1}.$
For a given $\boldsymbol{h}\in A_{\kappa}\setminus\{\boldsymbol{0}\}$ we count the number of $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ such that $(1/N)\boldsymbol{h}\in\mathbb{X}(\boldsymbol{g},N)$. This is equal to the number of elements in the set
$\{\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}\ :\ \exists\ell\in\{0,1,\ldots,N-1\}\mbox{ such that }\boldsymbol{h}\equiv\ell\boldsymbol{g}\pmod{N}\}.$
For $\ell\not=0$ we have
$\boldsymbol{h}\equiv\ell\boldsymbol{g}\pmod{N}\ \mbox{ if and only if }\ \boldsymbol{g}\equiv\ell^{-1}\boldsymbol{h}\pmod{N},$
where $\ell^{-1}$ is the inverse of $\ell$ modulo $N$ (remember that $N$ is a prime). Hence, for every $\ell\not=0$ there is exactly one such $\boldsymbol{g}$. This means that the number of $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ such that $(1/N)\boldsymbol{h}\in\mathbb{X}(\boldsymbol{g},N)$ is at most $N$.
Therefore
|$\left|\left\{\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}\ :\ \kappa(\mathbb{X}(\boldsymbol{g},N))\leq\frac{\kappa}{N}\right\}\right|\leq N2^{d}\binom{\kappa+d}{d}.$
Like above, if
$N2^{d}\binom{\kappa+d}{d}<\frac{1-\delta}{3}N^{d},$ (17)
then the number of $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ such that $\kappa(\mathbb{X}(\boldsymbol{g},N))>\kappa/N$ is at least $((2+\delta)/3)N^{d}$. According to (15), condition (17) is satisfied if $32^{d}(\kappa+d)^{d}/d!<(1-\delta)N^{d-1}/$ 3, i.e., for $\kappa=\lceil 2^{-1}(d!\,(1-\delta)N^{d-1}/3)^{1/d}\rceil-d-1$. Hence, for at least $((2+\delta)/3)N^{d}$ many $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ we have
$\kappa(\mathbb{X}^{\bot}(\boldsymbol{g},N))\geq B_{d,\delta}^{\prime\prime}\frac{N^{(d-1)/d}}{N}=\frac{B_{d,\delta}^{\prime\prime}}{N^{1/d}}$ (18)
for some $B_{d,\delta}^{\prime\prime}>0$, depending only on $d$ and $\delta$.
Now we combine the three parts. We have:
  • for the set $A$ of all $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ which satisfy (13) we have $|A|\geq((2+\delta)/3)N^{d}$,
  • for the set $B$ of all $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ which satisfy (16) we have $|B|\geq((2+\delta)/3)N^{d}$,
  • for the set $C$ of all $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ which satisfy (18) we have $|C|\geq((2+\delta)/3)N^{d}$.
Now $A\cap B\cap C$ consists of all $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$, for which (13), (16) and (18) hold simultaneously and we have
$|A\cap B\cap C|=$ $N^{d}-|A^{c}\cup B^{c}\cup C^{c}|$
$\geq$ $N^{d}-|A^{c}|-|B^{c}|-|C^{c}|$
$=$ $N^{d}-3N^{d}+|A|+|B|+|C|$
$\geq$ $\delta N^{d}.$
Obviously, the number of elements is an integer and hence $|A\cap B\cap C|\geq\lceil\delta N^{d}\rceil$.
Finally, (16) and (18) lead to
$\frac{1}{\kappa(\mathbb{X}(\boldsymbol{g},N))\kappa(\mathbb{X}^{\bot}(\boldsymbol{g},N))}\leq\frac{1}{B_{d,\delta}^{\prime}B_{d,\delta}^{\prime\prime}}.$
The estimate on $\rho_{p}(P(\boldsymbol{g},N))$ follows with (12). ∎
Let $\mathbb{P}$ be the set of prime numbers. Since the sequence of prime numbers $(p_{n})_{n\geq 1}$, where $p_{n}$ is the $n$-th prime number, satisfies $p_{n+1}<2p_{n}$ by Bertrand’s postulate, we obtain that there exists a sequence of rank-1 lattice point sets $(P_{N})_{N\in\mathbb{P}}$ which are quasi-uniform and which have low star-discrepancy.
Remark 3.8.
The result of Theorem 3.7 can be generalized to any lower-dimensional projection. That is, for given $d\in\mathbb{N}$, $p\in[1,\infty]$, and $\delta\in(0,1)$ for any $s\in\{1,\ldots,d\}$, there exist quantities $B_{s,p,\delta},C_{s,p,\delta}>0$ such that, for fixed non-empty $\mathfrak{u}\subseteq\{1,\ldots,d\}$ and for every prime number $N$ the bounds
$\rho_{p}(P_{\mathfrak{u}}(\boldsymbol{g},N))\leq B_{|\mathfrak{u}|,p,\delta}\qquad\mbox{and}\qquad D_{N}^{\ast}(P_{\mathfrak{u}}(\boldsymbol{g},N))\leq C_{|\mathfrak{u}|,p,\delta}\frac{(\log N)^{|\mathfrak{u}|}}{N},$ (19)
hold for at least $\lceil\delta N^{d}\rceil$ vectors $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$, where $P_{\mathfrak{u}}(\boldsymbol{g},N)\subseteq[0,1)^{|\mathfrak{u}|}$ denotes the projection of the point set $P(\boldsymbol{g},N)$ onto the coordinates indexed by $\mathfrak{u}$.
This implies that for any $\delta\in(1-2^{-d},1)$, there exists at least one vector $\boldsymbol{g}\in\{0,1,\ldots,N-1\}^{d}$ such that (19) holds simultaneously for all $\mathfrak{u}\subseteq\{1,\ldots,d\}$, i.e., for all lower-dimensional projections of the lattice point set. In this case, the quantities $B_{|\mathfrak{u}|,p,\delta}$ and $C_{|\mathfrak{u}|,p,\delta}$ depend both on the dimension $d$ through their dependence on $\delta$. Thus there exists a sequence of lattice point sets $(P(\boldsymbol{g},N))_{N\in\mathbb{P}}$, where $\mathbb{P}$ is the set of prime numbers, for which all projections onto the coordinates in $\mathfrak{u}\subseteq\{1,\ldots,d\}$ are quasi-uniform and for which all projections have small discrepancy.
3.3 $(n\boldsymbol{\alpha})$-sequences
So-called $(n\boldsymbol{\alpha})$-sequences (sometimes also refered to as Kronecker-sequences) are sequences of the form ($(\{n\boldsymbol{\alpha}\})_{n\geq 1}$, where $\boldsymbol{\alpha}=(\alpha_{1},\ldots,\alpha_{d})\in\mathbb{R}^{d}$ (sometimes in literature the sequence starts with $\boldsymbol{0}=0\boldsymbol{\alpha}$). These are important and well-studied objects in uniform distribution- and discrepancy-theory (see, e.g, [11, Section 1.4] and the references therein). It follows easily from Weyl’s criterion that an $(n\boldsymbol{\alpha})$-sequence is uniformly distributed modulo 1 if and only if $1,\alpha_{1},\ldots,\alpha_{d}$ are linearly independent over the rationals. Usually, results on the discrepancy of $(n\boldsymbol{\alpha})$-sequences depend on Diophantine properties of the real vector $\boldsymbol{\alpha}$. Regarding this topic a multitude of results are available in the literature, see, e.g., [11, 23, 27] and the references therein.
In this section we characterize those $\boldsymbol{\alpha}$, for which the $(n\boldsymbol{\alpha})$-sequences have the property of being quasi-uniform. The one-dimensional case has already been discussed in [16].
For $x\in\mathbb{R}$ let $\langle x\rangle=\min_{k\in\mathbb{Z}}|x-k|$ denote the distance of $x$ to the nearest integer. For a vector $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{d})\in\mathbb{R}^{d}$ let $\langle\boldsymbol{x}\rangle=\max_{1\leq j\leq d}\langle x_{j}\rangle$.
Theorem 3.9.
For $\boldsymbol{\alpha}\in\mathbb{R}^{d}$ the $(n\boldsymbol{\alpha})$-sequence ($(\{n\boldsymbol{\alpha}\})_{n\geq 1}$ is quasi-uniform, if and only if there exists a constant $c>0$ such that
$\langle n\boldsymbol{\alpha}\rangle\geq\frac{c}{n^{1/d}}\qquad\mbox{for all }n\in\mathbb{N}.$ (20)
Remark 3.10.
Property (20) characterizes so-called badly approximable vectors $\boldsymbol{\alpha}$, which is an important concept in Diophantine approximation theory.
Proof of Theorem 3.9.
First we show that (20) is sufficient for quasi-uniformity. The bound on the covering radius follows for instance directly from [25, Theorem 1]. This can be understood in the following way: The assumption in [25, Theorem 1] is stated in terms of an extremal form. However, by the Khintchine transference principle (see [5, Chapter V, Section 3, Proof of Theorem III]), (20) is equivalent to the assertion that the form $n_{1}\alpha_{1}+\cdots+n_{d}\alpha_{d}-n$ is extremal, that is,
$\langle n_{1}\alpha_{1}+\cdots+n_{d}\alpha_{d}\rangle\geq\frac{c^{\prime}}{(\max_{1\leq j\leq d}|n_{j}|)^{d}}\quad\mbox{for all }(n_{1},\ldots,n_{d})\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\}.$
Hence [25, Theorem 1] applies, thus we obtain a bound on the isotropic discrepancy (see Secton 1.2) of order $N^{-1/d}$. Since Euclidian balls are convex, the volume of the largest empty ball, which is exactly the covering radius, is bounded by the isotropic discrepancy, which itself is under the assumption of (20) of order of magnitude $N^{-1/d}$.
Now consider separation. Let $N\geq M\geq 1$. Then (20) implies
$\langle\{N\boldsymbol{\alpha}\}-\{M\boldsymbol{\alpha}\}\rangle=\langle(N-M)\boldsymbol{\alpha}\rangle\geq\frac{c}{(N-M)^{1/d}}\geq\frac{c}{N^{1/d}}.$
Thus we obtain
$\|\{N\boldsymbol{\alpha}\}-\{M\boldsymbol{\alpha}\}\|_{\infty}\geq\frac{c}{N^{1/d}}.$
Summing up, (20) implies quasi-uniformity.
In order to prove the converse direction, assume that (20) does not hold. Then for every $\varepsilon>0$ there exists an $n_{0}$ with
$\langle n_{0}\boldsymbol{\alpha}\rangle\leq\frac{\varepsilon}{n_{0}^{1/d}}.$
Consider two points $\boldsymbol{v}=(v_{1},\dots,v_{d}):=\{n_{0}\boldsymbol{\alpha}\}$ and $\boldsymbol{u}=(u_{1},\dots,u_{d}):=\{2n_{0}\boldsymbol{\alpha}\}$. Then for any $j\in\{1,\ldots,d\}$ we have $u_{j}=2v_{j}$ if $20\leq v_{j}<1/$ 2 and $u_{j}=2v_{j}-1$ otherwise, and thus we have
$\langle u_{j}-v_{j}\rangle=\langle v_{j}\rangle\quad\text{and}\quad|u_{j}-v_{j}|=\left\{\begin{array}[]{ll}|v_{j}|&\text{if $0\leq v_{j}<1/2$}\\ |1-v_{j}|&\text{otherwise}\end{array}\right\}=\langle v_{j}\rangle.$
Thus we have
$\|\boldsymbol{u}-\boldsymbol{v}\|_{\infty}=\langle\boldsymbol{v}\rangle\leq\frac{\varepsilon}{n_{0}^{1/d}}.$ (21)
If the $(n\boldsymbol{\alpha})$-sequence is quasi-uniform, then there exists a $c_{d}>0$ such that
$\|\boldsymbol{u}-\boldsymbol{v}\|_{\infty}=\|\{2n_{0}\boldsymbol{\alpha}\}-\{n_{0}\boldsymbol{\alpha}\}\|_{\infty}\geq\frac{c_{d}}{(2n_{0})^{1/d}}.$
However, in (21) the $\varepsilon$ can be chosen arbitrarily close to 0, and this yields a contradiction. Thus, the $(n\boldsymbol{\alpha})$-sequence cannot be quasi-uniform and the result follows.
From [5, Chapter 1; Proof of Theorem VIII] it follows that vectors $\boldsymbol{\alpha}$ exist which satisfy Condition (20). In particular, see [5, Chapter 1; Eq. (9)]. Theorem 3.9 generalizes the result shown in [16] for one-dimensional Kronecker sequences. Also there, it was shown that the sufficient and necessary condition for the sequence to be quasi-uniform is that the real number $\alpha$ is badly approximable. This condition is different from the condition for the sequence to be of low discrepancy. In the multi-dimensional case, [25, Theorem 1] shows that the isotropic discrepancy $J_{N}(\mathcal{S})$ of $(n\boldsymbol{\alpha})$-sequences $\mathcal{S}$ satisfying (20) is bounded by $J_{N}(\mathcal{S})\ll N^{-1/d}$ and the same upper bound applies to the star-discrepancy of $\mathcal{S}$. In particular, Condition (20) implies that the star-discrepancy $D_{N}^{\ast}$ of an $(n\boldsymbol{\alpha})$-sequence tends to 0 as $N$ goes to infinity and hence we obtain the following result:
Corollary 3.11.
If an $(n\boldsymbol{\alpha})$-sequence is quasi-uniform, then it is uniformly distributed.
The result of Corollary 3.11 is a specific feature of $(n\boldsymbol{\alpha})$-sequences which must not necessarily hold for general sequences. A counterexample is the Sobol’ sequence, which is uniformly distributed (its discrepancy is of order of magnitude $(\log N)^{d}/N$) but which is not quasi-uniform in dimension $2$, as shown in [17].
The bounds on the star-discrepancy mentioned above are not optimal. But we can provide explicit examples of $(n\boldsymbol{\alpha})$-sequences which are quasi-uniform and have at the same time a very low (in fact, almost best possible) star-discrepancy.
We will use the following discrepancy bound for Kronecker-sequences (see, e.g, [27, Theorem 6.1], [8, Section 4.8] or [23, Chapter 2, Section 3, Exercises 3.17]):
Let $\alpha_{1},\ldots,\alpha_{d}$ be irrational numbers. If for all $\varepsilon>0$ there is a $c(\varepsilon)>0$ such that
$(\overline{n}_{1}\cdots\overline{n}_{d})^{1+\varepsilon}\langle n_{1}\alpha_{1}+\cdots+n_{d}\alpha_{d}\rangle>c(\varepsilon)\quad\mbox{for all }(n_{1},\ldots,n_{d})\in\mathbb{Z}^{d}\setminus\{\boldsymbol{0}\},$ (22)
where $\overline{n}:=\max(1,|n|)$, then for the discrepancy of the corresponding $(n\boldsymbol{\alpha})$-sequence $\mathcal{S}$ we have
$D_{N}^{\ast}(\mathcal{S})\ll\frac{1}{N^{1-\delta}}\quad\mbox{for all }\delta>0.$
It was shown by Schmidt [39] that (22) holds whenever $\alpha_{1},\ldots,\alpha_{d}$ are algebraic and such that $1,\alpha_{1},\ldots,\alpha_{d}$ are linearly independent over $\mathbb{Q}$ (this generalizes the famous Thue-Siegel-Roth Theorem).
Let now $1,\alpha_{1},\ldots,\alpha_{d}$ be linear independent elements (hence a basis) of an algebraic extension of $\mathbb{Q}$ of degree $d+1$. Then, according to [5, Chapter V, Theorem 3], $\alpha_{1},\ldots,\alpha_{d}$ also satisfy property (20). Hence the corresponding $(n\boldsymbol{\alpha})$-sequence $\mathcal{S}$ in this case is quasi-uniform and its star-discrepancy satisfies
$D_{N}^{\ast}(\mathcal{S})\ll\frac{1}{N^{1-\delta}}\quad\mbox{for all }\delta>0.$
As a concrete example we can choose for instance
$\boldsymbol{\alpha}=(2^{1/(d+1)},2^{2/(d+1)},\ldots,2^{d/(d+1)})\in\mathbb{R}^{d}.$
Then the corresponding $(n\boldsymbol{\alpha})$-sequence $\mathcal{S}$ is quasi-uniform and its star-discrepancy is of order $N^{-1+\delta}$ for any $\delta>0$.